OpenStudy (anonymous):

At what point on the graph y=1/2x^2 is the tangent line parallel to the line?

OpenStudy (anonymous):

what line is the question referring to?

OpenStudy (anonymous):

sorry, i missed that. It's 2x-4y=3

OpenStudy (anonymous):

ok... so all we need is the slope of the line 2x - 4y = 3 so we can set the slope of the tangent line equal to that. did you find the derivative of y = 1/2x^2 ?

OpenStudy (anonymous):

x?

OpenStudy (anonymous):

is that the derivative?

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

ok.... since the derivative gives the slope of the tangent line at any point, then all we need to do now is set that to the slope of the line 2x - 4y = 3. what's the slope of 2x - 4y = 3 ????

OpenStudy (anonymous):

y=1/2x - 3/4

OpenStudy (anonymous):

just the slope.... what is it?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

ok... good... lemme summarize....

OpenStudy (anonymous):

dy/dx = x slope = 1/2 so if we set dy/dx = slope, we have: x = 1/2 this is the x-coordinate of the point on the graph where the tangent line is parallel to 2x - 4y = 3. do you know how to get the y-coordinate?

OpenStudy (anonymous):

Is that where we look for the x value?

OpenStudy (anonymous):

no... we already have the x value. we only need the y value (coordinate) of the point on the graph.

OpenStudy (anonymous):

hint: how do you plot points on the graph of y = 1/2x^2 ? how do you calculate the y-value when x=1/2 ?

OpenStudy (anonymous):

need another hint?